# RS-ISRefiner: Towards Better Adapting Vision Foundation Models for Interactive Segmentation of Remote Sensing Images

### v1

@@ Abstract
Interactive image segmentation(IIS) plays a critical role in generating precise annotations for remote sensing imagery, where objects often exhibit scale variations, irregular boundaries and complex backgrounds. However, existing IIS methods, primarily designed for natural images, struggle to generalize to remote sensing domains due to limited annotated data and computational overhead. To address these challenges, we proposed RS-ISRefiner, a novel click-based IIS framework tailored for remote sensing images. The framework employs an adapter-based tuning strategy that preserves the general representations of Vision Foundation Models while enabling efficient learning of remote sensing–specific spatial and boundary characteristics. A hybrid attention mechanism integrating convolutional local modeling with Transformer-based global reasoning enhances robustness against scale diversity and scene complexity. Furthermore, an improved probability map modulation scheme effectively incorporates historical user interactions, yielding more stable iterative refinement and higher boundary fidelity. Comprehensive experiments on six remote sensing datasets, including iSAID, ISPRS Potsdam, SandBar, NWPU, LoveDA Urban and WHUBuilding, demonstrate that RS-ISRefiner consistently outperforms state-of-the-art IIS methods in terms of segmentation accuracy, efficiency and interaction cost. These results confirm the effectiveness and generalizability of our framework, making it highly suitable for high-quality instance segmentation in practical remote sensing scenarios.
@@ I Introduction
Interactive Image Segmentation(IIS) is an effective approach that enables high-precision, pixel-level annotation with minimal human interaction. By continuously incorporating user feedback during the segmentations such as bounding boxes[1], clicks[2][3] and scribbles[4], IIS can accurately delineate object boundaries and has been widely applied in tasks including medical image analysis, natural scene understanding and remote sensing image interpretation. Compared with fully automatic segmentation methods, IIS offers higher controllability and flexibility, demonstrating significant advantages when handling ambiguous or fine-grained targets. For example, in remote sensing imagery, IIS can effectively distinguish shadow-coverd buildings from adjacent vegetation or identify irregularly shaped sandbars within rivers, with only a small amount of user interaction required to correct model predition errors.
In recent years, the emergence of Vision Foundation Models (VFMs) has ushered in new opportunities for the field of Interactive Image Segmentation (IIS). Vision Foundation Models (VFMs) are typically built upon large-scale pre-trained Convolutional Neural Networks (CNNs) and Transformers (e.g., ResNet[36], ConvNeXt[10], ViT[11], and Swin Transformer[12]), possessing rich transferable representations learned from massive and diverse datasets. Currently, in the field of IIS, numerous studies such as BRS[3], RITM[2] and SimpleClick[5] adopt VFMs as feature extraction encoders, and design additional downstream decoders to complete feature parsing and segmentation mask generation, as showed in Fig.1(a). However, most existing VFMs serving as feature encoders are pre-trained exclusively on large-scale natural images, and their performance degrades significantly when directly applied to remote sensing images without further training. For instance, although SAM[8] and its variants demonstrate superior performance on natrual images, they require numerous corrective clicks to handle RS scenarios such as small vehicles segmentation(sensitive to scale variation) or shadow-coverd buildings(with noisy boundaries), leading to substantially increased interaction costs[9]. To address this issue, most existing IIS methods perform task-specific full-parameter fine-tuning on the encoders of VFMs for interactive segmentation of remote sensing targets. Nevertheless, training VFMs consumes extensive computational resources, and full-parameter fine-tuning may overwrite the general visual knowledge acquired during the pre-training phase, thereby reducing the models’ inherent generalization capability[6]. Additionally, annotated data in the remote sensing domain is scarce and costly to obtain, which limits the improvement of model generalization in this specific field[7].
@@
With the advancement of multimodal large models, several fine-tuning adaptation algorithms (e.g., Adapter[13] and Prompt[14]) have also evolved. By leveraging fine-tuning mechanisms, pre-trained backbone networks can be efficiently transferred to downstream tasks, reducing training costs while preserving general visual knowledge[15]. For interactive segmentation in remote sensing, leveraging VFMs enables the preservation of general visual priors while selectively optimizing high-level features to better align with the spatial characteristic of RS imagery, as showed in Fig.1(b). For instance, adapter modules can be employed to capture domain-specific edge structures of RS targets such as the linear boundaries of building rooftops or the elongated contours of roads, without altering the parameters of the pretrained backbone network[16].
@@
However, existing IIS framework still face two major limitations when applied to remote sensing imagery. First, most methods(e.g., SimpleClick[5] and AdaptiveClick[17]) fail to fully exploit the historical information of user interactions. Although some approaches(e.g., MFP[18]) attempt to propagate such information through probability map modualtion, conventional modualtion mechanisms with fiexed circular modualtion windows are not well-suited to the irregular shapes of remote sensing targets. As a result, prediction instability or oscillation often occurs during multi-round interactions, particularly when dealing with geometrically complex objects such as river sandbars. Second, although VFMs provide powerful visual priors, directly transferring them to remote sensing imagery often leads to degraded boundary accuracy due to domain discrepancies. The visual features learned from natural images (e.g., object textures and edge details) do not align well with the spectral and spatial characteristics of remote sensing objects (e.g., spectral reflectance and spatial configuration)[19]. Moreover, existing adaptation modules such as the Adapter in SAM-HQ[20] are not specifically designed for remote sensing scenarios, making it difficult to capture the intrinsic spectral–spatial correlations of ground objects[6]. Therefore, there is an urgent need for an IIS framework that integrates an efficient VFM-based adaptation mechanism tailored for remote sensing with a robust historical information propagation strategy, enabling high-precision segmentation with minimal user interaction cost.
@@
In response to the above issues, we proposed a novel click-based IIS framework for remote sensing images, incorporating the probability map modulation mechanism and the backbone-freezing adapter fine-tuning algorithm to address the unique challenges of remote sensing data. Drawing on the insights of MFP[18], we leverage probability map modulation to propagate historical interaction information, thereby enhancing the model’s capability to retain object shapes and spatial contexts during iterative processes. With respect to the design of the adapter fine-tuning algorithm, we introduce one module to capture local spatial correlations and edge features specific to remote sensing objects, eliminating the need to retrain the entire backbone network.
@@ II-A Interactive Image Segmentation
Figure 2: Overall Structure. Our architecture consists of three main components: a frozen Vision Foundation Model (VFM) backbone (blue region), a parameter-trained extractor and encoder (orange-red region), and a probability map modualtion module at the top of the diagram. Both the training and testing processes follow a recursive workflow.
@@ II-B Vision Foundation Models and Parameter-efficient Fine-Tuning
Vision Foundation Models have become the driving force of modern computer vision. Early convolutional neural networks (CNNs), such as FCN[34], VGG[35], ResNet[36], DenseNet[37] and HRNet[38], provided effective feature extraction for pixel-level tasks, while the introduction of Transformers[39], such as ViT[11] and Swin Transformer[12], significantly enhanced global modeling and muti-scale representation. More recently, with the advent of large-scale pretraining, models including MAE[40], SAM[8], DINOv2[41] and ConvNeXt[10] have demonstrated remarkable generalization across tasks and domains. These models have achieved state-of-the-art performance in classification, detection and segmentation, establishing themselves as universal feature extractors with strong potential for remote sensing and interactive segmentation[42][43].
Nevertheless, despite their representational power, directly applying VFMs to interactive segmentation in remote sensing remains challenging due to high computational costs and the limited availability of annotated data. To bridge this gap, a range of parameter-efficient fine-tuning (PEFT) strategies have been developed[44]. Representative PEFT approaches include Adapter Tuning[13], Prompt Tuning[14] and LoRA[45], which reduce training overhead by introducing trainable modules while keeping the pretrained backbone largely frozen. Prompt Tuning requires minimal parameters but often offers limited flexibility for complex visual tasks. LoRA achieves high efficiency and incurs no inference overhead[45], but its low-rank formulation can restrict representational capacity in domains with large distribution gaps such as natural images to RS images. In contrast, Adapter Tuning strikes a favorable balance between efficiency and expressiveness, enabling tasks-specific adaptation with relatively small parameter costs while maintaining robust performance[13].
@@ II-C Historical Information Utilization in Interactive Refinement
Apart from architectural adaptation, IIS performance heavily depends on how effectively a model leverages historical user interactions during iterative refinement. RITM[2] and f-BRS[25] pioneered the use of previous segmentation masks as network input in the current segmentation round. The two models showed that even a simple feed-forward model using previous masks can enhance model stability without additional optimization. Since then, recent interactive segmentation methods have all adopted previous segmentation masks as network input. FocalClick[26] and FocusCut[29] refined results within a local window by correcting global predictions with local ones around clicks. GPCIS[49] formulated the problem of interactive segmentation as a Gaussian process classification task. EMC-Click[28] used self-attention and correlation modules to spread click information to unclicked areas. MST[32] addressed the issue of target scale variations in interactive segmentation by selectively fusing important multi-scale tokens based on user clicks to enhance the distinguishability between target and background tokens. Although the above methods all use previous masks or previous probability maps in the current round, none of them attempt to explicitly extract effective information from the probability maps. MFP[18] uses gamma correction-based probability map modulation to amplify probabilities near positive clicks and suppress those near negative clicks within a local window, and introduces late fusion of probability-related and backbone features to let historical context influence final prediction. However, the modulation mechanism of MFP[18] struggles with irregularly bounded remote sensing targets, and needs to be improved.
@@ III-A Overall Structure
To effectively adapt large-scale Vision Foundation Models (VFMs) to the remote sensing interactive segmentation task, our method adopts a modular backbone–adapter architecture that enables efficient fine-tuning while preserving the pretrained model’s generalization capability. As illustrated in Fig.2, the interactive segmentation model proposed in this paper mainly consists of four components: a frozen VFM Backbone, a High-Quality Feature Extractor, a Transformer Decoder and a Probability Map Modulation Module. The framework follows a recursive segmentation-refinement workflow, in which the model iteratively updates its prediction according to user interactions, thereby achieving progressively refined segmentation results with minimal clicks.
@@
Since this paper adopts the idea of probability map modulation from MFP[18], in each interaction round, the input to the network here consists of the image $I\in\mathbb{R}^{3\times H\times W}$, the previous segmentation mask $M_{prev.}\in\mathbb{R}^{1\times H\times W}$, the previous modulated segmentation mask $M_{mod.}\in\mathbb{R}^{1\times H\times W}$ and the click maps $M_{click}\in\mathbb{R}^{2\times H\times W}$ generated from the positive and negative clicks set. In the case of the initial click, since $M_{prev.}$ and $M_{mod.}$ do not exist, it is necessary to set both masks as all-zero matrices. Furthermore, $M_{prev.}$, $M_{mod.}$ and $M_{click}$ are concatenated into a unified tensor $M\in\mathbb{R}^{4\times H\times W}$ and jointly processed with the image $I$ through the frozen VFM Backbone to obtain stable and semantically rich representations. The High-Quality Feature Extractor then complements these backbone features with fine-grained spatial and boundary information, which are fused through cross-attention operations to generate enhanced local-global feature maps. Consistent with the methodology employed in MFP[18], $M_{prev.}$, $M_{mod.}$ and $I$ are subjected to Conv. and XceptionConv. operations to extract the corresponding probability map feature, which are subsequently fed into the decoder. This paper retains this approach, with the sole modification being that in the input components, image $I$ is substituted with the click map $M_{click}$, as illustrated in Fig.3. Research findings from [31] have demonstrated that encoding the click map and transmitting it to the decoder contributes to enhancing the model’s localization performance. The three inputs are processed by the Prev-Mask Processing Module, ultimately outputting the mask feature $F_{mask}$. Subsequently, the Transformer Decoder integrates image features, mask features and a learnable mask token to produce an updated instance mask $M_{seg.}$ that balances contextual consistency with edge precision. Finally, the predicted mask undergoes probability map modulation to correct local inconsistencies based on the spatial distribution of user clicks, yielding an updated probability map $M_{mod.}$ that is propagated to the next iteration.
@@
Figure 3: Previous Mask Processing Module. The module is mainly composed of four convolutional modules and one Xception convolutional module, where the light orange region at the bottom shows the detailed structure of the Xception convolutional module.
@@ III-B Frozen VFM Backbone
where $W$ is the weight matrix in the linear transformation. $b$ is the bias term in the linear transformation. $\sigma$ denotes the Sigmoid activation function, which serves to map the result of the linear transformation to the interval $\left[0,1\right]$, thereby obtaining the gating weights. The fused feature $F_{early}$ captures mid-level contextual and spatial details, complementing the high-level semantic features extracted by deeper layers. To address the scale variation inherent in remote sensing imagery, the final ViT block output $F_{ViT}^{N}$ is fed into the SimpleFPN to generate multi-scale representations $F_{1}^{ms},F_{2}^{ms},F_{3}^{ms},F_{4}^{ms}$}, which are fused into a unified multi-scale feature map $F_{multiscale}$. The multi-scale feature will be fed into the Transformer Decoder, enabling the network to maintain robustness across diverse object sizes.

### v2

# VFM-ISRefiner: Towards Better Adapting Vision Foundation Models for Interactive Segmentation of Remote Sensing Images

@@ Abstract
Interactive image segmentation(IIS) plays a critical role in generating precise annotations for remote sensing imagery, where objects often exhibit scale variations, irregular boundaries and complex backgrounds. However, existing IIS methods, primarily designed for natural images, struggle to generalize to remote sensing domains due to limited annotated data and computational overhead. To address these challenges, we propose VFM-ISRefiner, a novel click-based IIS framework tailored for remote sensing images. The framework employs an adapter-based tuning strategy that preserves the general representations of Vision Foundation Models while enabling efficient learning of remote-sensing–specific spatial and boundary characteristics. A hybrid attention mechanism integrating convolutional local modeling with Transformer-based global reasoning enhances robustness against scale diversity and scene complexity. Furthermore, an improved probability map modulation scheme effectively incorporates historical user interactions, yielding more stable iterative refinement and higher boundary accuracy. Comprehensive experiments on six remote sensing datasets, including iSAID, ISPRS Potsdam, SandBar, NWPU, LoveDA Urban and WHUBuilding, demonstrate that VFM-ISRefiner consistently outperforms state-of-the-art IIS methods in terms of segmentation accuracy, efficiency and interaction cost. These results confirm the effectiveness and generalizability of our framework, making it highly suitable for high-quality instance segmentation in practical remote sensing scenarios. The codes are available at https://github.com/wondelyan/VFM-ISRefiner .
@@ I Introduction
Interactive Image Segmentation(IIS) is an effective approach that enables high-precision, pixel-level annotation with minimal human interaction. By continuously incorporating user feedback during the segmentations such as bounding boxes[1], clicks[2][3] and scribbles[4], IIS can accurately delineate object boundaries and has been widely applied in tasks including medical image analysis, natural scene understanding and remote sensing image interpretation. Compared with fully automatic segmentation methods, IIS offers higher controllability and flexibility, demonstrating significant advantages when handling ambiguous or fine-grained targets[5]. For example, in remote sensing imagery, IIS can effectively distinguish shadow-covered buildings from adjacent vegetation or identify irregularly shaped sandbars within rivers, with only a small amount of user interaction required to correct model prediction errors.
In recent years, the emergence of Vision Foundation Models (VFMs) has ushered in new opportunities for the field of Interactive Image Segmentation (IIS). Vision Foundation Models (VFMs) are typically built upon large-scale pretrained Convolutional Neural Networks (CNNs) and Transformers (e.g., ResNet[6], ConvNeXt[7], ViT[8], and Swin Transformer[9]), possessing rich transferable representations learned from massive and diverse datasets[10]. Currently, in the field of IIS, numerous studies such as BRS[3], RITM[2] and SimpleClick[11] adopt VFMs as feature extraction encoders, and design additional downstream decoders to complete feature parsing and segmentation mask generation, as showed in Fig.1(a). However, most existing VFMs serving as feature encoders are pretrained exclusively on large-scale natural images, and their performance degrades significantly when directly applied to remote sensing images without further training. For instance, although SAM[12] and its variants demonstrate superior performance on natural images, they require numerous corrective clicks to handle RS scenarios such as small vehicles segmentation(sensitive to scale variation) or shadow-covered buildings(with noisy boundaries), leading to substantially increased interaction costs[13]. To address this issue, most existing IIS methods perform task-specific full-parameter fine-tuning on the encoders of VFMs for interactive segmentation of remote sensing targets. Nevertheless, training VFMs consumes extensive computational resources, and full-parameter fine-tuning may overwrite the general visual knowledge acquired during the pre-training phase, thereby reducing the models’ inherent generalization capability[14]. Additionally, annotated data in the remote sensing domain is scarce and costly to obtain[15], which limits the improvement of model’s generalization in this specific domain[16].
@@
With the advancement of multimodal large models, several fine-tuning adaptation algorithms (e.g., Adapter[17] and Prompt[18]) have also evolved. By leveraging fine-tuning mechanisms, pretrained backbone networks can be efficiently transferred to downstream tasks, reducing training costs while preserving general visual knowledge[19]. For interactive segmentation in remote sensing, leveraging VFMs enables the preservation of general visual priors while selectively optimizing high-level features to better align with the spatial characteristic of RS imagery, as showed in Fig.1(b). For instance, adapter modules can be employed to capture domain-specific edge structures of RS targets such as the linear boundaries of building rooftops or the elongated contours of roads, without altering the parameters of the pretrained backbone network[20].
@@
However, existing IIS frameworks still face two major limitations when applied to remote sensing imagery. First, most methods(e.g., SimpleClick[11] and AdaptiveClick[21]) fail to fully exploit the historical information of user interactions. Although some approaches(e.g., MFP[22]) attempt to propagate such information through probability map modulation, conventional modulation mechanisms with fixed circular modulation windows are not well-suited to the irregular shapes of remote sensing targets. As a result, prediction instability or oscillation often occurs during multi-round interactions, particularly when dealing with geometrically complex objects such as river sandbars. Second, although VFMs provide powerful visual priors, directly transferring them to remote sensing imagery often leads to degraded boundary accuracy due to domain discrepancies. The visual features learned from natural images (e.g., object textures and edge details) do not align well with the spectral and spatial characteristics of remote sensing objects (e.g., spectral reflectance and spatial configuration)[23]. Moreover, existing adaptation modules such as the Adapter in SAM-HQ[24] are not specifically designed for remote sensing scenarios, making it difficult to capture the intrinsic spectral-spatial correlations of ground objects[14]. Therefore, there is an urgent need for an IIS framework that integrates an efficient VFM-based adaptation mechanism tailored for remote sensing with a robust historical information propagation strategy, enabling high-precision segmentation with minimal user interaction cost.
@@
In response to the above issues, we proposed a novel click-based IIS framework for remote sensing images: VFM-ISRefiner, incorporating the probability map modulation mechanism and the backbone-freezing adapter fine-tuning algorithm to address the unique challenges of remote sensing data. Drawing on the insights of MFP[22], we leverage probability map modulation to propagate historical interaction information, thereby enhancing the model’s capability to retain object shapes and spatial contexts during iterative processes. With respect to the design of the adapter fine-tuning algorithm, we introduce a module to capture local spatial correlations and edge features specific to remote sensing objects, eliminating the need to retrain the entire backbone network.
@@ II-B Vision Foundation Models and Parameter-efficient Fine-Tuning
Vision Foundation Models have become the driving force of modern computer vision. Early convolutional neural networks (CNNs), such as FCN[38], VGG[39], ResNet[6], DenseNet[40] and HRNet[41], provided effective feature extraction for pixel-level tasks, while the introduction of Transformers[42], such as ViT[8] and Swin Transformer[9], significantly enhanced global modeling and multi-scale representation. More recently, with the advent of large-scale pretraining, models including MAE[43], SAM[12], DINOv2[44] and ConvNeXt[7] have demonstrated remarkable generalization across tasks and domains. These models have achieved state-of-the-art performance in classification, detection and segmentation, establishing themselves as universal feature extractors with strong potential for remote sensing and interactive segmentation[45][46].
Nevertheless, despite their representational power, directly applying VFMs to interactive segmentation in remote sensing remains challenging due to high computational costs. To bridge this gap, a range of parameter-efficient fine-tuning (PEFT) strategies have been developed[47]. Representative PEFT approaches include Adapter Tuning[17], Prompt Tuning[18] and LoRA[48], which reduce training overhead by introducing trainable modules while keeping the pretrained backbone largely frozen. Prompt Tuning requires minimal parameters but often offers limited flexibility for complex visual tasks. LoRA achieves high efficiency and incurs no inference overhead[48], but its low-rank formulation can restrict representational capacity in domains with large distribution gaps such as natural images to RS images. In contrast, Adapter Tuning strikes a favorable balance between efficiency and expressiveness, enabling tasks-specific adaptation with relatively small parameter costs while maintaining robust performance[17].
@@ II-C Historical Information Utilization in Interactive Refinement
Apart from architectural adaptation, IIS performance heavily depends on how effectively a model leverages user’s historical interactions during iterative refinement. RITM[2] and f-BRS[29] pioneered the use of previous segmentation masks as network input in the current segmentation round. The two models showed that even a simple feed-forward model using previous masks can enhance model stability without additional optimization. Since then, interactive segmentation methods have all adopted previous segmentation masks as network input. FocalClick[30] and FocusCut[33] refined results within a local window by correcting global predictions with local ones around clicks. GPCIS[52] formulated the problem of interactive segmentation as a Gaussian process classification task. EMC-Click[32] used self-attention and correlation modules to spread click information to unclicked areas. MST[36] addressed the issue of target scale variations in interactive segmentation by selectively fusing important multi-scale tokens based on user clicks to enhance the distinguishability between target and background tokens. Although the above methods all use previous masks or previous probability maps in the current round, none of them attempt to extract more effective information from the probability maps. MFP[22] uses gamma correction-based probability map modulation to amplify probabilities near positive clicks and suppress those near negative clicks within a local window, and introduces late fusion of probability-related and backbone features to let historical context influence final prediction. However, the modulation mechanism of MFP[22] struggles with irregularly bounded remote sensing targets, and needs to be improved.
Figure 2: Overall Structure. Our architecture consists of three main components: a frozen Vision Foundation Model (VFM) backbone (blue region), a parameter-trained extractor and decoder (orange-red region) and a probability map modulation module at the top of the diagram. Both the training and testing processes follow a recursive workflow.
@@ III-A Overall Structure
To effectively adapt large-scale Vision Foundation Models (VFMs) to remote sensing interactive segmentation tasks, our method adopts a modular backbone-adapter architecture that enables efficient fine-tuning while preserving the pretrained model’s generalization capability. As illustrated in Fig.2, the interactive segmentation model proposed in this paper mainly consists of four components: a frozen VFM Backbone, a High-Quality Feature Extractor, a Transformer Decoder and a Probability Map Modulation Module. The framework follows a recursive segmentation-refinement workflow, in which the model iteratively updates its prediction according to user interactions, thereby achieving progressively refined segmentation results with minimal clicks.
@@
Since this paper adopts the idea of probability map modulation from MFP[22], in each interaction round, as shown in Fig.2(a), the input to the network here consists of the image $I\in\mathbb{R}^{3\times H\times W}$, the previous segmentation mask $M_{prev.}\in\mathbb{R}^{1\times H\times W}$, the previous modulated segmentation mask $M_{mod.}\in\mathbb{R}^{1\times H\times W}$ and the click maps $M_{click}\in\mathbb{R}^{2\times H\times W}$ generated from the positive and negative clicks set. In the case of the initial click, since $M_{prev.}$ and $M_{mod.}$ do not exist, it is necessary to set both masks as all-zero matrices. Furthermore, as illustrated in Fig.2(b), $M_{prev.}$, $M_{mod.}$ and $M_{click}$ are concatenated into a unified tensor $M\in\mathbb{R}^{4\times H\times W}$ and jointly processed with the image $I$ through the frozen VFM Backbone to obtain stable and semantically rich representations. The High-Quality Feature Extractor presented in Fig.2(c) then complements these backbone features with fine-grained spatial and boundary information, which are fused through cross-attention operations to generate enhanced local-global feature maps. Consistent with the methodology employed in MFP[22], $M_{prev.}$, $M_{mod.}$ and $I$ are subjected to Conv. and XceptionConv. operations to extract the corresponding probability map feature, which are subsequently fed into the decoder. This paper retains this approach, with the sole modification being that in the input components, image $I$ is substituted with the click map $M_{click}$, as illustrated in Fig.3. Research findings from [35] have demonstrated that encoding the click map and transmitting it to the decoder contributes to enhancing the model’s localization performance. The three inputs are processed by the Prev-Mask Processing Module, ultimately outputting the mask feature $F_{mask}$. Subsequently, the Transformer Decoder in Fig.2(d) integrates image features, mask features and a learnable mask token to produce an updated instance mask $M_{seg.}$ that balances contextual consistency with edge precision. Finally, as shown in Fig.2(e) the predicted mask undergoes probability map modulation to correct local inconsistencies based on the spatial distribution of user clicks, yielding an updated probability map $M_{mod.}$ that is propagated to the next iteration.
@@
Figure 3: Previous Mask Processing Module. The module is mainly composed of four convolutional modules and one Xception-convolutional module, where the light orange region at the bottom shows the detailed structure of the Xception-convolutional module.
@@ III-B Frozen VFM Backbone
where $W$ is the weight matrix in the linear transformation. $b$ is the bias term in the linear transformation. $\sigma$ denotes the Sigmoid activation function, which serves to map the result of the linear transformation to the interval $\left[0,1\right]$, thereby obtaining the gating weights. The fused feature $F_{early}$ captures mid-level contextual and spatial details, complementing the high-level semantic features extracted by deeper layers. To address the scale variation inherent in remote sensing imagery, the final ViT block output $F_{ViT}^{N}$ is fed into the SimpleFPN to generate multi-scale representations $F_{1}^{ms},F_{2}^{ms},F_{3}^{ms},F_{4}^{ms}$}, which are fused into a unified multi-scale feature map $F_{multiscale}$. The multi-scale feature will be fed into the trainable transformer decoder, enabling the network to maintain robustness across diverse object sizes.